# EquiBot: SIM(3)-Equivariant Diffusion Policy for Generalizable and Data Efficient Learning

### v1

@@ B.1 Failure Analysis
B.2 Additional Real Robot Results
@@
Detailed performance analysis of the laundry door closing task. To understand the performance of diffusion policy [1] and our method better, we perform a more detailed analysis in the Laundry Door Closing task. In Table 2, we report four different metrics of policy performance in each evaluation setup. Success Rate measures the percentage of evaluations that end within the success criteria we set; Close with Click Sound measures the percentage of episodes that end with the laundry door closed completely after making a clicking sound; Total Accumulated Missing Angle measures the sum of the opening angles of the laundry door at the end of the 10 evaluation episodes; Collision or Safety Issue measures the percentage of evaluation runs that are terminated because of undesired collisions or critical safety issues, such as the robot arm getting stuck at the laundry door. From the evaluation results, we see that the DP policy not only has a lower success rate but also has a much larger accumulated missing angle. The baseline also suffers from many safety issues requiring episodes to be manually terminated by the robot operator. Our method has much fewer safety issues when it executes.
Qualitative results. In Figure 9, we show qualitative rollout samples for all evaluation scenarios we mentioned in the main paper, plus one bonus task where two robots lift a woven basket onto a coffee table.
@@
Supplementary Materials C Equivariant Distributions and Diffusion
In this section, we show that with equivariance in the per-step diffusion update, the final output action distribution is also equivariant under stochasticity.
We mainly discuss equivariance to $\mathrm{SO}(3)$-rotations in the diffusion process. Equivariance to translations and scaling (to get SIM(3)-equivariant architectures) is achieved via canonicalization before the diffusion process.
@@
In our case, we have observation ${\mathbf{O}}_{t}$ as the condition and the diffusion process of the actions ${\mathbf{A}}_{t}^{0:K}$. The prior distribution is a standard Gaussian distribution $p({\mathbf{A}}_{t}^{K}|{\mathbf{O}}_{t})$, which is equivariant to $\mathrm{SO}(3)$ transformations. The transition probabilities $p({\mathbf{A}}_{t}^{k-1}|{\mathbf{A}}_{t}^{k},{\mathbf{O}}_{t})$ are predictions by an equivariant network, and thus are equivariant to $\mathrm{SO}(3)$-rotations. Therefore, the final action output $p({\mathbf{A}}_{t}^{0}|{\mathbf{O}}_{t})$ is also $\mathrm{SO}(3)$-equivariant.
Supplementary Materials D Network Architectures and Implementation Details
In this section, we describe in detail the architecture of our method (Figure 10 gives a visualization).
Observation and action spaces. In all simulated and real robot tasks except for Push T, we use 13-dimensional proprioception information and a 7-dimensional action space for each robot. The proprioception data for each robot consists of the following information: a 3-dimensional end-effector position, a 6-dimensional vector denoting end-effector orientation (represented by two columns of the end-effector rotation matrix), a 3-dimensional vector indicating the direction of gravity, and a scalar that represents the degree to which the gripper is opened. The action space for each robot consists of the following information: a 3-dimensional vector for the end-effector position velocity, a 3-dimensional vector for the end-effector angular velocity in axis-angle format, and a scalar denoting the gripper action.
@@
Encoder architecture. In all tasks except for Push T, we use a SIM(3)-equivariant version of PointNet++ with 4 layers and hidden dimensionality 128. In the Push T task, we decrease the number of layers to 2, since the number of points in the point cloud observation is much smaller in this task.
Figure 10: EquiBot Architecture. Given input scene point cloud, robot proprioception, noisy actions, and the diffusion timestamp, our architecture processes position, direction, and scalar information independently, uses the encoder outputs to scale position information into position and scale invariant values, and then routes them into an SO(3)-equivariant conditional U-net to predict denoised actions. In the figure, we omit scaling for the ease of viewing. VecLinear and VecConv1D refer to a SO(3)-equivariant version of linear and convolution 1D layers.
@@
Supplementary Materials E Real Robot Setup and Experiments – Further Details
E.1 Human Demo Parsing Infrastructure
We use a single ZED 2 camera to record human natural motion in real-time, which is more flexible and time-efficient than the expert demonstration from robot teleoperation. With that, we create a set of human demonstrations $\mathcal{D}=\{\tau^{n}\}_{n=1}^{N}$. Each human demonstration consists of a series of RGB-D image frames $\tau^{n}=\{I_{t}^{n}\}_{t=1}^{T}$, where $T$ is the episode horizon.
@@ E.1 Human Demo Parsing Infrastructure
E.2 Mobile Robot Control Infrastructure
@@ E.3 Task Details
Supplementary Materials F Simulation Tasks – Further Details
@@
Push T. In this task, a 2D anchor pushes a T shaped object on a plane of dimension $512\times 512$ pixels. The task reward is computed as the percentage of the T shape that overlaps with the target T pose.

### v2

@@ C.1 Failure Analysis
C.2 Additional Real Robot Results
@@
Detailed performance analysis of the laundry door closing task. To understand the performance of the Diffusion Policy [8] and our method better, we perform a more detailed analysis in the laundry door closing task. In Table 2, we report four different metrics of policy performance in each evaluation setup. Success Rate measures the percentage of evaluations that end within the success criteria we set; Close with Click Sound measures the percentage of episodes that end with the laundry door closed completely after making a clicking sound; Total Accumulated Missing Angle measures the sum of the opening angles of the laundry door at the end of the 10 evaluation episodes; Collision or Safety Issue measures the percentage of evaluation runs that are terminated because of undesired collisions or critical safety issues, such as the robot arm getting stuck at the laundry door. From the evaluation results, we see that the DP policy not only has a lower success rate but also has a much larger accumulated missing angle. The baseline also suffers from many safety issues requiring episodes to be manually terminated by the robot operator. Our method has much fewer safety issues when it executes.
Qualitative results. In Figure 12, we show qualitative rollout samples for all evaluation scenarios we mentioned in the paper, plus one bonus task where two robots lift a woven basket onto a coffee table.
@@
Appendix D Equivariant Distributions and Diffusion
In this section, we prove Proposition 1 that we introduced in Section 3.2 of the main paper.
Proposition 1.
Let $p(x^{K}|c)$ be a SO(3)-equivariant density function conditioned on $c$, i.e. $\forall{\mathbf{R}}\in SO(3),p(x^{K}|c)=p({\mathbf{R}}x^{K}|{\mathbf{R}}c)$. If Markov transitions $p(x^{k-1}|x^{k},c)$ are SO(3)-equivariant for all $k$, i.e. $p(x^{k-1}|x^{k},c)=p({\mathbf{R}}x^{k-1}|{\mathbf{R}}x^{k},{\mathbf{R}}c)$, then the density $p(x^{0}|c)=\int p(x^{T}|c)\Pi_{k=1}^{K}p(x^{k-1}|x^{k},c)$ is also SO(3)-equivariant.
Proof.
@@ Appendix D Equivariant Distributions and Diffusion
Appendix E Method Architectures and Implementation Detail
In this section, we describe in detail the architecture of our method. We visualize the architecture of our model in Figure 13.
Observation and action spaces. In all simulated and real robot tasks except for Push T, we use a 13-dimensional proprioception information and a 7-dimensional action space for each robot. The proprioception data for each robot consists of the following information: a 3-dimensional end-effector position, a 6-dimensional vector denoting end-effector orientation (represented by two columns of the end-effector rotation matrix), a 3-dimensional vector indicating the direction of gravity, and a scalar that represents the degree to which the gripper is opened. The action space for each robot consists of the following information: a 3-dimensional vector for the end-effector position velocity, a 3-dimensional vector for the end-effector angular velocity in axis-angle format, and a scalar denoting the gripper action.
@@ Appendix E Method Architectures and Implementation Detail
Encoder architecture. In all tasks except for Push T, we use a SIM(3)-equivariant version of PointNet++ with 4 layers and hidden dimensionality 128. In the Push T task, we decrease the number of layers to 2 since the number of points in the point cloud observation is much smaller in this task.
Figure 13: Architecture of EquiBot. Given input scene point cloud, robot proprioception, noisy actions, and the diffusion timestamp, our architecture processes position, direction, and scalar information independently, uses the encoder outputs to scale position information into position and scale invariant values, and then routes them into an SO(3)-equivariant conditional U-net to predict denoised actions. In the figure, we omit scaling for the ease of viewing. VecLinear and VecConv1D refer to a SO(3)-equivariant version of linear and convolution 1D layers.
@@
Normalization. Data normalization can be important to the performance of diffusion models. Vanilla diffusion policy normalizes the observations and actions separately. We instead normalize all 3D-vector inputs (including observation and action) together due to our SIM(3)-equivariance assumptions. Implementation-wise, we take a subset of training data and compute the mean point cloud scale and mean action scale as $s_{\text{pc}}$ and $s_{\text{ac}}$. Then, all position- and velocity-related information is divided by $s_{\text{pc}}/s_{\text{ac}}$ at the start of the network forward pass and multiplied back before the output is returned. This normalization factor ensures that the diffusion process always works with actions with values within the $-1$ to $1$ range. We normalize scalar information in the same way as the vanilla diffusion policy.
Appendix F Real Robot Setup and Experiments – Further Details
F.1 Human Demo Parsing Infrastructure
We use a single ZED 2 camera to record human natural motion in real-time, which is way more flexible and time-efficient than the expert demonstration from human teleoperation of a robot. With that, we create a set of human demonstrations $\mathcal{D}=\{\tau^{n}\}_{n=1}^{N}$. Each human demonstration consists of a series of RGB-D image frames $\tau^{n}=\{I_{t}^{n}\}_{t=1}^{T}$, where $T$ is the episode horizon.
@@ F.1 Human Demo Parsing Infrastructure
F.2 Mobile Robot Control Infrastructure
@@ F.3 Task Details
Appendix G Simulation Tasks – Further Details
@@
Push T. In this task, a 2D anchor pushes a T-shaped object on a plane of dimension $512\times 512$ pixels. The task reward is computed as the percentage of the T shape that overlaps with the target T pose.
@@
Appendix H Limitations and Future Work – Further Details